ABC Corporation currently has an Inventory Turnover of 11.59, an Accounts Receivable Turnover of 19.69. How many days are in the operating cycle?

Enter your answer rounded off to two decimal points.
Answer #1
Inventory turnover = COGS/inventory
Inventory turnover = 29097/10853
Inventory turnover = 2.68
days of inventory on hand = number of days in a year/inventory turnover
days of inventory on hand = 365/11.59
days of inventory on hand = 31.49
Receivables turnover = Credit sales/receivables
Receivables turnover = 146625/35971
Receivables turnover = 4.08
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/19.69
days of sales outstanding = 18.54
Operating cycle = days of sales outstanding + days of inventory on hand
Operating cycle = 18.54+31.49
Operating cycle = 50.03